OpenStudy (anonymous):

solve the equation x^2+2x=2

OpenStudy (saifoo.khan):

decimals.

OpenStudy (saifoo.khan):

x = 0.73 x = -2.73

myininaya (myininaya):

x^2+2x+1=2+1 (x+1)^2=3 x+1=plus or minus sqrt{3} x=-1 plus or minus sqrt{3} x=-1+sqrt{3} or x=-1-sqrt{3}